What effect would chlorine (Cl2), hypochlorite (OCl-), and nitrite (NO2-) have on the dissolved oxygen of a water sample? Write equations to represent their individual interactions.
1
Expert's answer
2020-11-20T10:02:49-0500

When they react with water they all form a weak acid

Cl2 + H2O = HOCl + H+ + Cl-

OCl- + H2O = HOCl + OH-

NO2- + H2O = HNO2 + H+